Five forces 2, 4, 6, 8 and 10 N, respectively, at the same point, from angles of 72 degrees with each other. Find the resultant.




1
Expert's answer
2021-02-22T10:24:30-0500
"R_x=2+4\cos{\frac{2\pi}{5}}+6\cos{\frac{4\pi}{5}}+8\cos{\frac{6\pi}{5}}+10\cos{\frac{8\pi}{5}}\\R_x=-5"

"R_y=4\sin{\frac{2\pi}{5}}+6\sin{\frac{4\pi}{5}}+8\sin{\frac{6\pi}{5}}+10\sin{\frac{8\pi}{5}}\\R_y=-6.88"

"R=8.5"